Solve the following system of equations. x + 2y = 2 - 4x + 3y = 25
Answer #1

Sogn Given System of equations x+2y = 2 -4x+3y =25 1 2 let, A= [ ; B= -4 3 i.e 2 -4 3 + [as] :x=[] | as] [ :/:~***** 6:31]<-4